OpenStudy (sleepyjess):

When using the equation y^2 = 4ax to find the equation of a parabola, how do I know what "a" is?

OpenStudy (sleepyjess):

Example, Question: Find an equation of the parabola with vertex at (0, 0) and focus at (3, 0). Solution: The distance form the vertex (0, 0) to the focus (3, 0) is a = 3. Based on equation 2, the equation of this parabola is \(y^2 = 4ax\) \(y^2 = 12x\)

OpenStudy (sleepyjess):

Where do they get the 3 from?

TheSmartOne (thesmartone):

They get the 3 from the focus. http://prntscr.com/9ds9no

TheSmartOne (thesmartone):



OpenStudy (sleepyjess):

This was a huge struggle point last year for me >_<

OpenStudy (sleepyjess):

So the x coordinate of the focus is "a"?

TheSmartOne (thesmartone):

Yup :)

OpenStudy (sleepyjess):

So, Question: Find an equation of the parabola with vertex at (1, 2) and focus at (4, 8). Would "a" still be 4?

TheSmartOne (thesmartone):

And depending on the equation, the graph will look differently. And from the equation you can tell where the focus would be by knowing what 'a' is :)

OpenStudy (sleepyjess):

Idek if that would be a possible vertex and focus, just an example

TheSmartOne (thesmartone):

For that kind of question, you have to use another formula: (x-h)^2 = 4p(y-k)

OpenStudy (sleepyjess):

How do I know what h, k and p are?

OpenStudy (sleepyjess):

Vertex is (h, k) right?

TheSmartOne (thesmartone):

As always, (h,k) is the vertex

TheSmartOne (thesmartone):

And, for p "The absolute value of p is the distance between the vertex and the focus and the distance between the vertex and the directrix."

OpenStudy (sleepyjess):

My book says 4a, is there any difference between p and a?

TheSmartOne (thesmartone):

I guess not lol

TheSmartOne (thesmartone):

http://www.purplemath.com/modules/parabola3.htm :)